OpenStudy (anonymous):

Using the replacement set, find the solutions for the equation. y = 0.5x + 10 Replacement Set: {(-2, 9.5), (0, 10),(3, 11.5),(4, 13)}

OpenStudy (kropot72):

You just need to plug the x and y values in each pair into the equation and find which ones make the equation true. For example using (-2, 9.5): \[9.5=(0.5\times-2)+10=-1+10=9\] Obviously the first pair is not a solution. Can you try the second pair in the set?

OpenStudy (anonymous):

Thanks that really helps

OpenStudy (anonymous):

Algebra is so not my thing

OpenStudy (anonymous):

Trying to help my kiddo

OpenStudy (kropot72):

To test the second pair, put 10 in place of y and 0 in place of x.

OpenStudy (anonymous):

These are his options for answers so the 9 throws me for a loop {(-2, 9.5), (0, 10),(3, 11.5),(4, 13)} {(-2, 9.5), (0, 10),(3, 11.5)} {(0, 10),(3, 11.5),(4, 13)} { (0, 10),(3, 11.5)}

OpenStudy (kropot72):

I already showed you that (-2, 9.5) is not a solution, so the first two choices are not correct.

OpenStudy (anonymous):

oh duh sorry

OpenStudy (anonymous):

lol

OpenStudy (kropot72):

I understand your question. Obviously 9.5 does not equal 9. Can you test the second pair of values?

OpenStudy (anonymous):

i think it's the last one

OpenStudy (kropot72):

Yes, the last choice is correct.

OpenStudy (anonymous):